# TimeMosaic: Temporal Heterogeneity Guided Time Series Forecasting via Adaptive Granularity Patch and Segment-wise Decoding

### v2

@@ Abstract
Codehttps://github.com/Day333/TimeMosaic
@@ 1 Instruction
Multivariate time series forecasting (Yu et al. 2024, 2023; Zheng et al. 2023) plays a critical role in numerous real-world domains, such as finance (Zhu et al. 2024a, b), transportation (Shu, Cai, and Xiong 2021), climate (Schneider and Dickinson 1974), healthcare (Morid, Sheng, and Dunbar 2023), and energy management (Liu et al. 2023). These applications require accurate and efficient modeling of intricate temporal relationships among multiple correlated variables. Recently, forecasting methods employing patch-based representations have shown remarkable performance and have become increasingly prevalent (Chen et al. 2024; Jin et al. 2024; Kudrat et al. 2025; Nie et al. 2023). Patch-based approaches excel at capturing localized temporal structures and mitigating noise through segmented encoding. Recently proposed time series models, such as Sundial (Liu et al. 2025), also use patch-based time series segmentation to organize event contexts for benchmarking time series foundation models. Typically, these methods divide input sequences into fixed-length patches, implicitly assuming uniform information density and temporal complexity across the entire sequence.
@@
However, real-world time series often exhibit significant variability in local information density (Huang et al. 2023; Warren Liao 2005). Segments characterized by complex or abrupt changes inherently contain higher information densities, whereas smoother or stationary regions possess relatively lower densities. Current methods employing fixed-size temporal segments disregard this inherent variability, resulting in inadequate modeling of information-rich regions and redundant encoding in smoother ones. Beyond local variability, empirical evidence (Xie et al. 2025) also reveals two key structural properties: motif reuse, characterized by Zipf-like frequency distributions, and structural clarity, measured by latent-space separability (see Section 2 for definitions and illustrations), which further highlight a structural limitation of fixed-length patching: it cannot simultaneously preserve reusable long-range patterns and maintain well-delimited local boundaries. A similar limitation has been observed in autoregressive image generation, such as DQ-VAE (Huang et al. 2023), where fixed-size region encodings likewise fail to adapt to diverse structural patterns.
@@
To address these two challenges, we propose TimeMosaic, a novel forecasting framework that integrates adaptive patch embedding with segment-wise decoding. On the input side, we design an adaptive granularity patch embedding strategy inspired by dynamic quantization (Huang et al. 2023). Our approach segments input sequences into variable-length patches based on local temporal information density, balancing motif reuse with structural clarity. Unlike prior multi-granularity methods that produce overlapping or re-ordered patches, TimeMosaic ensures that each time step belongs to exactly one patch, thereby preserving strict temporal continuity that is essential for forecasting. As illustrated in Figure 1, fixed-length patching overlooks temporal heterogeneity: it loses details in information-dense regions and introduces redundancy in smooth intervals.
On the output side, we introduce a segment-wise prediction framework based on multi-task prompt tuning (Liu et al. 2021; Crawshaw 2020a). By treating different horizons as distinct yet related subtasks, our method equips the model with segment-aware prompt embeddings that capture horizon-specific difficulty and information requirements without modifying the backbone parameters. This design enables specialization across prediction segments, effectively leveraging the rich and heterogeneous information captured by adaptive patches. As illustrated in Figure 1, TimeMosaic adopting segment-wise forecasting that treats each horizon as a dedicated subtask.
@@ 2 Related Work
Recent advances in time series forecasting have been driven by patch-based representations. PatchTST (Nie et al. 2023) first introduced temporal patches into Transformer-based models, followed by Patchwise (Kudrat et al. 2025), which utilized local statistics, and TimeFilter (Hu et al. 2025), which applied fine-grained filtration. Time-LLM (Jin et al. 2024) further demonstrated the utility of patches in LLM-based forecasting. However, all these models adopt fixed-length patching, assuming uniform temporal complexity and lacking flexibility to adapt to local variations.
To overcome this limitation, multi-granularity patching has been explored. PatchMLP (Tang and Zhang 2025) introduced multiple segment lengths to enrich temporal features, while PathFormer (Chen et al. 2024) and DualSG (Ding et al. 2025) selected patch sizes based on importance scores or maps. Although effective, these methods may break temporal consistency due to overlapping or misordered patches. For instance, coarse patches can appear earlier than fine ones in the sequence, distorting input chronology and undermining forecasting coherence, see Appendix Figure  7.
@@
Recent research (Xie et al. 2025) shows that temporal patches in time series follow a Zipf-like frequency distribution (Axtell 2001), exhibit clear clustering properties (measured by Silhouette scores (Rousseeuw 1987)), and display strong sequential dependencies. Zipf conformity reveals the emergence of reusable temporal motifs, enabling compact representation and pattern reuse, while structural clarity reflects how distinctly different temporal patterns can be separated in latent space, facilitating stable representation learning and reliable generalization.
@@
Autoregressive (AR) models (Lin et al. 2023; Khaldi et al. 2023; Stankeviciute, M Alaa, and Van der Schaar 2021) such as RNNs and LSTMs (Kong et al. 2025; Yadav and Thakkar 2024) have long dominated time series forecasting. Even recent foundation models (Shi et al. 2025; Liu et al. 2024b; Ansari et al. 2024) largely follow the AR formulation. However, these methods suffer from exposure bias (Ranzato et al. 2016) and equal treatment of short- and long-term predictions, despite their differing complexity and uncertainty.
@@
Multi-task learning (MTL) (Ruder 2017; Crawshaw 2020b) enhances generalization by sharing knowledge across tasks. In the context of large language models, it has been applied to summarization (Laban et al. 2023), QA (Zhuang et al. 2023), and sentiment analysis (Zhang et al. 2024; Cao et al. 2024), often in combination with parameter-efficient methods like LoRA (Xia et al. 2024) and prompt tuning (Lester, Al-Rfou, and Constant 2021).
We follow this trend by introducing prompt-based MTL into time series forecasting. Compared to basis-reconstruction methods such as TimeBase (Huang et al. 2025), which rely on strong periodicity assumptions and static segment structures, our model learns dynamic, segment-aware prompts for flexible and efficient temporal adaptation. This enables better modeling of local heterogeneity without assuming low-rank structure or modifying the shared encoder.
Figure 3: Overall architecture of TimeMosaic, which shares the same modular arrangement with the encoder of Transformer. (a) The input multivariate time series is first processed by the Adaptive Patch Embedding module, which segments the sequence into patches of varying granularity based on learnable region-aware decisions. (b) A set of learnable prompt tokens are injected into the input sequence and interact with patch embeddings via multi-head attention to guide segment-wise forecasting. The attention maps in this panel visualize the interactions among adaptive patches as well as between patches and prompts. (c) Input sequence are normalized at the Channel level, same as iTransformer (Liu et al. 2024a). (d) The model performs segment-wise forecasting, from the short-term to the long-term.
@@ 4.3 Adaptive Patch Embedding
where $g_{r}$ is the predicted logits over candidate patch sizes for region $x_{r}$, and $\theta_{r}$ is the index of the selected patch size. The classifier is shared across all regions and channels, and is implemented as a two-layer MLP. To ensure end-to-end differentiability, we apply the Gumbel-Softmax (Jang, Gu, and Poole 2017) during training.
@@ 4.4 Prompt Tuning for Segment-wise Forecasting
We formulate multi-interval time series forecasting as a segment-wise multitask learning problem, where each forecast segment corresponds to an individual prediction subtask. To achieve parameter-efficient segment adaptation, we draw inspiration from Prompt Tuning (Lester, Al-Rfou, and Constant 2021; Liu et al. 2021) and extend it to the temporal domain. Specifically, we design learnable segment-aware prompts to inject segment-specific inductive biases into a shared encoder, allowing the model to specialize its attention behavior for each segment without modifying backbone parameters.
@@ 5 Experiments
We conduct extensive experiments on 17 real-world multivariate time series datasets, which are divided into long-term and short-term forecasting tasks. For long-term forecasting, we use ETTh1, ETTh2, ETTm1, ETTm2 (Zhou et al. 2021), Weather, Traffic, Electricity, ExchangeRate (Wu et al. 2023), Solar-Energy (Lai et al. 2018), and Wind (Location1–4) (Xie et al. 2025). For short-term forecasting, we adopt four benchmark traffic datasets: PEMS03, PEMS04, PEMS07, and PEMS08 (Wang et al. 2024, 2025b).
@@
We choose the last state-of-the-art LTSF models, including TimeFilter (Hu et al. 2025), SimpleTM (Chen et al. 2025), PatchMLP (Tang and Zhang 2025), xPatch (Stitsyuk and Choi 2025), DUET (Qiu et al. 2025), PathFormer (Chen et al. 2024), iTransformer (Liu et al. 2024a), TimeMixer (Wang et al. 2024), PatchTST (Nie et al. 2023), FreTS (Yi et al. 2023), DLinear (Zeng et al. 2023), LightTS (Zhang et al. 2022) as baselines for our experiments. We compared the ability of zero-shot with the GPT4TS (Zhou et al. 2023) and LLMTime (Nate Gruver and Wilson 2023).And we compare against pre-trained foundation time series models, including TimeMoE (Shi et al. 2025), MOIRAI (Woo et al. 2024), Chronos (Ansari et al. 2024), TimesFM (Das et al. 2024), and Moment (Goswami et al. 2024), using the parameter settings as specified in their respective publications.
@@
Although our core design centers on segment-wise forecasting with adaptive patch granularity, handling multivariate inputs remains essential. To ensure compatibility, we incorporate a modular channel modeling component, which supports both Channel-Independent (CI) (Nie et al. 2023) and Channel-Dependent (CD) (Hu et al. 2025) schemes. Specifically, we provide multiple instantiations ranging from simple per-variable modeling to prompt-augmented or calendar-aware variants. These options are fully pluggable and do not interfere with the main forecasting pipeline. By default, we adopt the Channel dependent strategy. A comprehensive description and comparison of all variants are provided in Appendix E.
@@
All the experiments are implemented in PyTorch (Paszke et al. 2019), and conducted on eight A800 GPU. Even in recent studies, it is common to set drop_last=Truedrop\_last=True during data preprocessing (Chen et al. 2025; Wang et al. 2025a), and this issue has been explicitly discussed in the TFB (Qiu et al. 2024), highlighting its potential impact on evaluation fairness. In our work, we adopt a unified benchmark with drop_last=Falsedrop\_last=False to ensure full sequence utilization and fair comparisons. Our benchmark comprehensively includes 20+ state-of-the-art forecasting models across 17 real-world datasets, ensuring a fair comparison by unifying implementation settings and avoiding over-tuning parameter. And some approaches (Hu et al. 2025) discard the use of EarlyStopping, which in our view disregards the essential role of the validation set. Therefore, when integrating with other methods, we consistently retain this setting to ensure fair and reliable evaluation.
@@
To ensure fair comparison, we conduct two types of experiments. For long-term forecasting, we follow the unified evaluation protocol proposed by TimesNet (Wu et al. 2023), using a lookback length $L=96$ and prediction lengths T=96,192,336,720T={96,192,336,720} across all long-term datasets. For short-term forecasting, we adopt a fixed prediction length $T=12$, which is commonly used in traffic prediction tasks. The lookback length is also set to $L=96$. The averaged results under this unified setting are reported in Table 1. Table 1 presents results across all models using best configurations obtained via hyperparameter tuning, including the number of layers, attention heads, hidden dimension ($d_{\text{model}}$), and feedforward dimension ($d_{\text{ff}}$), with a lookback window of 96. In contrast, Table 5 reports results with an identical parameter setting shared by all models, using a longer lookback window of 320.
@@
To assess cross-dataset generalization, we conduct zero-shot forecasting where models are trained on one dataset and tested on another. As shown in Appendix Table 23, TimeMosaic consistently achieves the best performance across all transfer settings, outperforming GPT4TS (Zhou et al. 2023), DLinear, and PatchTST. In addition, we further scale up TimeMosaic on the BLAST dataset (Shao et al. 2025) with an input length of 512 and an output horizon of 720. This large-scale setting (27M parameters, trained on two V100 GPUs for about 40 hours) allows a direct comparison with recent time series foundation models (TSFMs). The results (Appendix K, Tables 4 and 16) show that TimeMosaic achieves zero-shot performance on par with representative TSFMs such as TimeMoe (Shi et al. 2025), Moirai (Woo et al. 2024), Chronos (Ansari et al. 2024), TimesFM (Das et al. 2024), and Moment (Goswami et al. 2024) while maintaining a moderate model size and inference cost. This confirms that the architectural choices of adaptive patch embedding and segment-wise decoding remain effective even at the foundation-model scale.
@@ Appendix B Experiment Details
In accordance with prior works such as (Wu et al. 2023), we adopt prediction lengths of $\{96,192,336,720\}$ for long-term forecasting tasks, and a fixed length of $12$ for short-term forecasting. However, we identify a common experimental flaw: the test data loader in these works often sets drop_last=True, which may discard up to batch_size - 1 samples during evaluation. This issue has been discussed in (Qiu et al. 2024), yet persists in recent literature (Chen et al. 2025; Wang et al. 2025a). To ensure fair comparison and full test coverage, we uniformly set drop_last=False in our experiments.

### v3

@@ Abstract
Codehttps://github.com/BenchCouncil/TimeMosaic
@@ 1 Instruction
Multivariate time series forecasting (yu2024ginar; DSformer; HybridZheng) plays a critical role in numerous real-world domains (qiu2025DBLoss; qiu2025dag), such as finance (lsrigru; mcigru), transportation (traffic_flow), climate (schneider1974climate), healthcare (health), and energy management (liu2023sadi). These applications require accurate and efficient modeling of intricate temporal relationships among multiple correlated variables. Recently, forecasting methods employing patch-based representations have shown remarkable performance and have become increasingly prevalent (chen2024pathformer; timellm; patchwise; patchtst). Patch-based approaches excel at capturing localized temporal structures and mitigating noise through segmented encoding. Recently proposed time series models, such as Sundial (liu2025sundial), also use patch-based time series segmentation to organize event contexts for benchmarking time series foundation models. Typically, these methods divide input sequences into fixed-length patches, implicitly assuming uniform information density and temporal complexity across the entire sequence.
@@
However, real-world time series often exhibit significant variability in local information density (Huang_2023_CVPR; WARRENLIAO20051857). Segments characterized by complex or abrupt changes inherently contain higher information densities, whereas smoother or stationary regions possess relatively lower densities. Current methods employing fixed-size temporal segments disregard this inherent variability, resulting in inadequate modeling of information-rich regions and redundant encoding in smoother ones. Beyond local variability, empirical evidence (xie2025languagetimelanguagemodel) also reveals two key structural properties: motif reuse, characterized by Zipf-like frequency distributions, and structural clarity, measured by latent-space separability (see Section 2 for definitions and illustrations), which further highlight a structural limitation of fixed-length patching: it cannot simultaneously preserve reusable long-range patterns and maintain well-delimited local boundaries. A similar limitation has been observed in autoregressive image generation, such as DQ-VAE (Huang_2023_CVPR), where fixed-size region encodings likewise fail to adapt to diverse structural patterns.
@@
To address these two challenges, we propose TimeMosaic, a novel forecasting framework that integrates adaptive patch embedding with segment-wise decoding. On the input side, we design an adaptive granularity patch embedding strategy inspired by dynamic quantization (Huang_2023_CVPR). Our approach segments input sequences into variable-length patches based on local temporal information density, balancing motif reuse with structural clarity. Unlike prior multi-granularity methods that produce overlapping or re-ordered patches, TimeMosaic ensures that each time step belongs to exactly one patch, thereby preserving strict temporal continuity that is essential for forecasting. As illustrated in Figure 1, fixed-length patching overlooks temporal heterogeneity: it loses details in information-dense regions and introduces redundancy in smooth intervals.
On the output side, we introduce a segment-wise prediction framework based on multi-task prompt tuning (liu2021gpt; crawshaw2020multitasklearningdeepneural). By treating different horizons as distinct yet related subtasks, our method equips the model with segment-aware prompt embeddings that capture horizon-specific difficulty and information requirements without modifying the backbone parameters. This design enables specialization across prediction segments, effectively leveraging the rich and heterogeneous information captured by adaptive patches. As illustrated in Figure 1, TimeMosaic adopting segment-wise forecasting that treats each horizon as a dedicated subtask.
@@ 2 Related Work
Recent advances in time series forecasting have been driven by patch-based representations. PatchTST (patchtst) first introduced temporal patches into Transformer-based models, followed by Patchwise (patchwise), which utilized local statistics, and TimeFilter (hu2025timefilter), which applied fine-grained filtration. Time-LLM (timellm) further demonstrated the utility of patches in LLM-based forecasting. However, all these models adopt fixed-length patching, assuming uniform temporal complexity and lacking flexibility to adapt to local variations.
To overcome this limitation, multi-granularity patching has been explored. PatchMLP (patchmlp) introduced multiple segment lengths to enrich temporal features, while PathFormer (chen2024pathformer) and DualSG (ding2025dualsg) selected patch sizes based on importance scores or maps. Although effective, these methods may break temporal consistency due to overlapping or misordered patches. For instance, coarse patches can appear earlier than fine ones in the sequence, distorting input chronology and undermining forecasting coherence, see Appendix Figure  7.
@@
Recent research (xie2025languagetimelanguagemodel) shows that temporal patches in time series follow a Zipf-like frequency distribution (zipf), exhibit clear clustering properties (measured by Silhouette scores (ROUSSEEUW198753)), and display strong sequential dependencies. Zipf conformity reveals the emergence of reusable temporal motifs, enabling compact representation and pattern reuse, while structural clarity reflects how distinctly different temporal patterns can be separated in latent space, facilitating stable representation learning and reliable generalization.
@@
Autoregressive (AR) models (lin2023segrnn; khaldi2023best; stankeviciute2021conformal) such as RNNs and LSTMs (kong2025unlockingpowerlstmlong; yadav2024noa) have long dominated time series forecasting. Even recent foundation models (shi2024timemoe; liu2024timer; ansari2024chronos) largely follow the AR formulation. However, these methods suffer from exposure bias (ranzato2016sequenceleveltrainingrecurrent) and equal treatment of short- and long-term predictions, despite their differing complexity and uncertainty.
@@
Multi-task learning (MTL) (ruder2017overviewmultitasklearningdeep; DBLP:journals/corr/abs-2009-09796) enhances generalization by sharing knowledge across tasks. In the context of large language models, it has been applied to summarization (laban-etal-2023-summedits), QA (NEURIPS2023_9cb2a749), and sentiment analysis (zhang-etal-2024-sentiment; cao2024enhanced), often in combination with parameter-efficient methods like LoRA (NEURIPS2024_747dc7c6) and prompt tuning (power).
We follow this trend by introducing prompt-based MTL into time series forecasting. Compared to basis-reconstruction methods such as TimeBase (huang2024timebase), which rely on strong periodicity assumptions and static segment structures, our model learns dynamic, segment-aware prompts for flexible and efficient temporal adaptation. This enables better modeling of local heterogeneity without assuming low-rank structure or modifying the shared encoder.
Figure 3: Overall architecture of TimeMosaic, which shares the same modular arrangement with the encoder of Transformer. (a) The input multivariate time series is first processed by the Adaptive Patch Embedding module, which segments the sequence into patches of varying granularity based on learnable region-aware decisions. (b) A set of learnable prompt tokens are injected into the input sequence and interact with patch embeddings via multi-head attention to guide segment-wise forecasting. The attention maps in this panel visualize the interactions among adaptive patches as well as between patches and prompts. (c) Input sequence are normalized at the Channel level, same as iTransformer (liu2023itransformer). (d) The model performs segment-wise forecasting, from the short-term to the long-term.
@@ 4.3 Adaptive Patch Embedding
where $g_{r}$ is the predicted logits over candidate patch sizes for region $x_{r}$, and $\theta_{r}$ is the index of the selected patch size. The classifier is shared across all regions and channels, and is implemented as a two-layer MLP. To ensure end-to-end differentiability, we apply the Gumbel-Softmax (jang2017categorical) during training.
@@ 4.4 Prompt Tuning for Segment-wise Forecasting
We formulate multi-interval time series forecasting as a segment-wise multitask learning problem, where each forecast segment corresponds to an individual prediction subtask. To achieve parameter-efficient segment adaptation, we draw inspiration from Prompt Tuning (power; liu2021gpt) and extend it to the temporal domain. Specifically, we design learnable segment-aware prompts to inject segment-specific inductive biases into a shared encoder, allowing the model to specialize its attention behavior for each segment without modifying backbone parameters.
@@ 5 Experiments
We conduct extensive experiments on 17 real-world multivariate time series datasets, which are divided into long-term and short-term forecasting tasks. For long-term forecasting, we use ETTh1, ETTh2, ETTm1, ETTm2 (haoyietal-informer-2021), Weather, Traffic, Electricity, ExchangeRate (timesnet), Solar-Energy (lai2018modelinglongshorttermtemporal), and Wind (Location1–4) (xie2025languagetimelanguagemodel). For short-term forecasting, we adopt four benchmark traffic datasets: PEMS03, PEMS04, PEMS07, and PEMS08 (wang2023timemixer; wang2024timemixer++).
@@
We choose the last state-of-the-art LTSF models, including TimeFilter (hu2025timefilter), SimpleTM (chen2025simpletm), PatchMLP (patchmlp), xPatch (stitsyuk2025xpatch), DUET (qiu2025duet), PathFormer (chen2024pathformer), iTransformer (liu2023itransformer), TimeMixer (wang2023timemixer), PatchTST (patchtst), FreTS (yi2023frequencydomain), DLinear (Zeng2022AreTE), LightTS (zhang2022morefastmultivariatetime) as baselines for our experiments. We compared the ability of zero-shot with the GPT4TS (zhou2023one) and LLMTime (gruver2023llmtime).And we compare against pre-trained foundation time series models, including TimeMoE (shi2024timemoe), MOIRAI (woo2024moirai), Chronos (ansari2024chronos), TimesFM (timesfm), and Moment (goswami2024moment), using the parameter settings as specified in their respective publications.
@@
Although our core design centers on segment-wise forecasting with adaptive patch granularity, handling multivariate inputs remains essential. To ensure compatibility, we incorporate a modular channel modeling component, which supports both Channel-Independent (CI) (patchtst) and Channel-Dependent (CD) (hu2025timefilter) schemes. Specifically, we provide multiple instantiations ranging from simple per-variable modeling to prompt-augmented or calendar-aware variants. These options are fully pluggable and do not interfere with the main forecasting pipeline. By default, we adopt the Channel dependent strategy. A comprehensive description and comparison of all variants are provided in Appendix E.
@@
All the experiments are implemented in PyTorch (paszke2019pytorchimperativestylehighperformance), and conducted on eight A800 GPU. Even in recent studies, it is common to set drop_last=Truedrop\_last=True during data preprocessing (chen2025simpletm; wang2025fredf), and this issue has been explicitly discussed in the TFB (qiu2024tfb), highlighting its potential impact on evaluation fairness. Following the time series forecasting benchmark TFB’s settings (qiu2024tfb), we do not use the ”drop last” trick during the testing phase to ensure a fair comparison. Our benchmark comprehensively includes 20+ state-of-the-art forecasting models across 17 real-world datasets, ensuring a fair comparison by unifying implementation settings and avoiding over-tuning parameter. And some approaches (hu2025timefilter) discard the use of EarlyStopping, which in our view disregards the essential role of the validation set. Therefore, when integrating with other methods, we consistently retain this setting to ensure fair and reliable evaluation.
@@
To ensure fair comparison, we conduct two types of experiments. For long-term forecasting, we follow the unified evaluation protocol proposed by TimesNet (timesnet), using a lookback length $L=96$ and prediction lengths T=96,192,336,720T={96,192,336,720} across all long-term datasets. For short-term forecasting, we adopt a fixed prediction length $T=12$, which is commonly used in traffic prediction tasks. The lookback length is also set to $L=96$. The averaged results under this unified setting are reported in Table 1. Table 1 presents results across all models using best configurations obtained via hyperparameter tuning, including the number of layers, attention heads, hidden dimension ($d_{\text{model}}$), and feedforward dimension ($d_{\text{ff}}$), with a lookback window of 96. In contrast, Table 5 reports results with an identical parameter setting shared by all models, using a longer lookback window of 320.
@@
To assess cross-dataset generalization, we conduct zero-shot forecasting where models are trained on one dataset and tested on another. As shown in Appendix Table 23, TimeMosaic consistently achieves the best performance across all transfer settings, outperforming GPT4TS (zhou2023one), DLinear, and PatchTST. In addition, we further scale up TimeMosaic on the BLAST dataset (blast) with an input length of 512 and an output horizon of 720. This large-scale setting (27M parameters, trained on two V100 GPUs for about 40 hours) allows a direct comparison with recent time series foundation models (TSFMs). The results (Appendix K, Tables 4 and 16) show that TimeMosaic achieves zero-shot performance on par with representative TSFMs such as TimeMoe (shi2024timemoe), Moirai (woo2024moirai), Chronos (ansari2024chronos), TimesFM (timesfm), and Moment (goswami2024moment) while maintaining a moderate model size and inference cost. This confirms that the architectural choices of adaptive patch embedding and segment-wise decoding remain effective even at the foundation-model scale.
@@ Appendix B Experiment Details
In accordance with prior works such as (timesnet), we adopt prediction lengths of $\{96,192,336,720\}$ for long-term forecasting tasks, and a fixed length of $12$ for short-term forecasting. However, we identify a common experimental flaw: the test data loader in these works often sets drop_last=True, which may discard up to batch_size - 1 samples during evaluation. This issue has been discussed in (qiu2024tfb), yet persists in recent literature (chen2025simpletm; wang2025fredf). To ensure fair comparison and full test coverage, we uniformly set drop_last=False in our experiments.